OpenStudy (anonymous):

PLEASE HELP ME I DESPERATELY NEED HELP ON THIS FOR MY GRADE- Express the product of 2x^2+6x-8 and x+3 in standard form?

OpenStudy (anonymous):

2x²+6x-8(x+3) 2x²*x=2x³ 2x²*3=6x² 6x*x=6x² 6x*3=18x -8*x=-8x -8*3=-24 6x²+6x²=12x² 18x-8x=10x 2x³+12x²+10x-24